OpenStudy (anonymous):

what is the factored form of (xy)^3-a^3?

OpenStudy (lgbasallote):

difference of two cubes a^3 - b^3 = (a-b)(a^2 + ab + b^2)

OpenStudy (cwrw238):

use the identity a^3 - b^3 = (a - b)(a^2 + ab + b^2)

OpenStudy (anonymous):

how (xy)^3-a^3 become trinomial?

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (anonymous):

if the (xy)^3-a^3 is the answer for x^3y^3-a^3.. what is the expanded form?

OpenStudy (lgbasallote):

use the formula i stated a^3 - b^3 = (a-b)(a^2 + ab + b^2)

OpenStudy (lgbasallote):

where are you getting confused?

OpenStudy (anonymous):

the expanded form is (a-b)(a^2 + ab + b^2). just substitute xy in for a and a in place of b.

OpenStudy (anonymous):

therefore it will be (xy - a) (xy^2 + xya +a^2)

OpenStudy (anonymous):

ohh , thank you :)

OpenStudy (lgbasallote):

careful there @michelle1503 remember your a is xy therefore a^2 will be (xy)^2 = x^2 y^2 not just xy^2 ;)

OpenStudy (anonymous):

pleasure. yeah my bad forgot the brackets. they make a huge difference just add them in. thanks for the heads up Igbasallote :)

OpenStudy (anonymous):

thank you both of you :)

OpenStudy (lgbasallote):

hehe no worries :) i usually have problems like that lots of timestoo

OpenStudy (anonymous):

pleasure :) oh my sole u have no idea how many silly mistakes i make on a daily basis haha.